OpenStudy (ttop0816):

Solve the following system of equations by the substitution method. 5x = y + 6 2x - 3y = 4 What is the value of the y-coordinate? -8/13 13/14 14/13

OpenStudy (mertsj):

Solve the first equation for y

OpenStudy (ttop0816):

using the second equation i got y= 2/3x-4/3!

OpenStudy (ttop0816):

then do i plug in that in to 5x= y(here!)+6?

OpenStudy (mertsj):

yes

OpenStudy (ttop0816):

i got 5x=2/3x-4/3 then would i bring the 2/3x over to the other side and then calculate?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

You forgot the +6

OpenStudy (mertsj):

\[5x=\frac{2}{3}x-\frac{4}{3}+6\]

OpenStudy (ttop0816):

then would i move 2/3x to the other side?? 5x-2/3x=4/3+6

OpenStudy (mertsj):

yes

OpenStudy (ttop0816):

then i got 13/3x= 14/3 ....?

OpenStudy (mertsj):

yes

OpenStudy (ttop0816):

after i would move 13/3 to the other side and make x=14/3-13/3 perhaps??

OpenStudy (mertsj):

No. Multiply both sides by the reciprocal of 13/3

OpenStudy (ttop0816):

then it would be x=14/13?

OpenStudy (mertsj):

yes. Now find y

OpenStudy (ttop0816):

it would be -8/13??

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

Now I would like to point out how much easier the work would have been if you would have solved the first equation for y instead of the second equation.